OpenStudy (anonymous):

Which of the following equations represents a line parallel to 4x - y = -5? Question 9 options: 8x - 2y = -10 4x + y = -2 4x - y = 2 2x - 1/2y = -5/2

OpenStudy (anonymous):

4x - y = -5 put it into slope-intercept form 4x - y = -5 -y = -5 - 4x y = 4x + 5 since y = mx + b we see that m = 4 and b = 5 for y1 = m1x + b1 and y2 = m2x + b2 to be parallel lines they must satisfy the condition that m1 = m2 and b1 =/= b2 (if the intercepts are equal as well its the same exact line!) so we need to see which option has a slope equal to 4 testing A: 8x - 2y = -10 -2y = -10 - 8x y = 4x + 5 while m = 4 which is the same as the given line B: 4x + y = -2 y = -2 - 4x m = -4 not 4 so B is disqualified C: 4x - y = 2 -y = 2 - 4x y = 4x - 2 m = 4 which is the same as the given line so this is parallel D: 2x - 1/2y = -5/2 -1/2 y = -5/2 - 2x -y = -5 - 4x y = 4x + 5 see the explanation for A answer is ultimately C buddy

OpenStudy (anonymous):

Thanx good explaining by the way ;)